OpenStudy (anonymous):

Use the distributive property to simplify: 25(2x + 1/5)

OpenStudy (vishweshshrimali5):

50x + 5

jhonyy9 (jhonyy9):

50x +25/5 = 50x +5 = 5(10x+1)